OpenStudy (anonymous):

what is 9.797958971132712 written as a fraction or sqrt96 answer as fraction

OpenStudy (anonymous):

im bad at changing decimals to fractions and vice versa

OpenStudy (ddcamp):

Hmmm... sqrt(96) = sqrt(6)×sqrt(16) = 4×sqrt(6). I don't know if that can easily be written as a fraction.

OpenStudy (anonymous):

on the problem it said any answers written as decimals will not be graded or accepted

OpenStudy (anonymous):

what if you just took the sqrt of 96 and wrote it as an improper fraction....

OpenStudy (ddcamp):

Did it say to simplify the radical? Because if that's what you're doing, you can leave some of the answer under the sqrt sign.

OpenStudy (anonymous):

it said to simplify it yeah \[\sqrt{96}\] i saw the fraction thing and kinda latched on to it my mistake lol

OpenStudy (ddcamp):

For example: To simplify sqrt(3×a²), you would say sqrt(3×a²) = sqrt(3) × sqrt(a²) = a×sqrt(3) Some of the answer is still under the radical, but it can't be simplified anymore.

OpenStudy (anonymous):

so how would i simplify a radical that doesnt have variables?

OpenStudy (unklerhaukus):

factorise the numbers into primes

OpenStudy (anonymous):

is the answer \[4\sqrt{6}\]

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

thanks everyone